OpenStudy (accidentalaichan):

if g(x) = x2 − 3, find g(4).

OpenStudy (anonymous):

plug in 4 for x

OpenStudy (anonymous):

g(x)=x^2+3 g(4)= 4^2+3=16+3 g(4)=19 (I believe)

OpenStudy (anonymous):

@Bella<3 you copied down the function incorrectly

OpenStudy (anonymous):

your work is correct though

OpenStudy (accidentalaichan):

I forgot to mention, I am looking for a step-by-step explanation ^^''

OpenStudy (anonymous):

Dang lol sorry lol

OpenStudy (calculusxy):

g(4)=4^2-3 16-3=13 13

OpenStudy (accidentalaichan):

@calculusxy, where did you get \[4^{2}\] ?

OpenStudy (anonymous):

under the assumption that x2 = x^2

OpenStudy (accidentalaichan):

Oh, I see.

OpenStudy (anonymous):

if you meant it as multiplication, then its easier to display the number before the variable like so 2x

OpenStudy (accidentalaichan):

Ah... I meant to be multiplying x by 2

OpenStudy (anonymous):

regardless, once you have made the substitution of x=4, you just follow order of operations otherwise known as pemdas or bodmas

OpenStudy (accidentalaichan):

This is my work: g(4) = 4 x 2 - 3 g(4)= 8 - 3 g(4) = 5 ? What am I doing wrong?

OpenStudy (anonymous):

nothing? your work looks fine as is

OpenStudy (accidentalaichan):

Oh, Okay Thanks so much!

OpenStudy (anonymous):

no problem

OpenStudy (calculusxy):

But aren't you supposed to square 4?

OpenStudy (calculusxy):

@AccidentalAiChan

OpenStudy (accidentalaichan):

No, x2 is x times 2, not x squared

OpenStudy (anonymous):

x2 is a completely valid method of writing 2 * x however it is often used as x^2 simply because people do not realize that the carrot "^" is used to denote exponents, which is why when people write down quadratic equations, it is often missing the ^ @calculusxy